OpenStudy (anonymous):

As a valet at the mall, Davis works 7-hour shifts. He earns $6 per hour and keeps 90% of his tip money. How much does Davis need to be tipped so that he takes home exactly $87.50, before taxes are deducted, per shift?

OpenStudy (anonymous):

Well first we have to figure out his base pay. Since Davis works 7 hours at 6$/hour, his base pay is: 7 x 6 = $42 so what does that mean? Well if Davis needs to make 87.50 and he already takes home 42: 87.50 - 42 = $45.50 is the amount of money he needs to be tipped. Here comes the tricky part. Since Davis only gets 90% of his tip money, 45.50 has to be 90% of a bigger number. We can set up a proportion and cross-multiply: 45.50/x = 90/100 90x = 4550 x = $55.555..... So Davis will need to make at least $55.56 in tips in order to take home 87.50 before taxes.

OpenStudy (anonymous):

Thanks your alot of help :)